OpenStudy (anonymous):

Let F(x) be the function (1/X+11) Then the quotient F(10+h)-F(10)/ h can be simplified to (-1/ah+b) for What is A= What is B = Can someone please show the steps, I got A right, but B was incorrect.

OpenStudy (anonymous):

\[f(x)=\frac{1}{x+11}\] \[f(10)=\frac{1}{22}\] \[f(10+h)-f(10)=\frac{1}{21+h}-\frac{1}{21}=\frac{21-21-h}{(21+h)(21)}=\frac{-h}{21(21+h)}\]

OpenStudy (anonymous):

divide by h and get \[\frac{-1}{21^2+21h}\]